OpenStudy (anonymous):

Evaluate the variable expression for x = 4 7 - 7 * (x - 5)

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

Failure to follow the guidelines may lead to your suspension or banning from the site. If you see any violations, please report them by using a report abuse link. OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution.

OpenStudy (anonymous):

7+7= 14

OpenStudy (anonymous):

yes ilbouskila

OpenStudy (anonymous):

the answer is 14

OpenStudy (anonymous):

7-7*(4-5) =7-7*-1 7+7= 14